OpenStudy (anonymous):

p+c+s=2543c+p=1503p=123+s what does c=

OpenStudy (lastdaywork):

Please write each equation in a separate line.

OpenStudy (anonymous):

p = 123, so c+ p(123) =1503 so subtract 123 then you get c, which is 1380

OpenStudy (texaschic101):

p + c + s = 2543 c + p = 1503 --> p = 1503 - c p = 123 + s (1503 - c) + c + s = 2543 1503 + s = 2543 s = 2543 - 1503 s = 1040 p = 123 + s p = 123 + 1040 p = 1163 1163 = 1503 - c 1163 - 1503 = - c -340 = -c 340 = c check... p + c + s = 2543 1163 + 340 + 1040 = 2543 2543 = 2543 (correct) c = 340

OpenStudy (texaschic101):

oh....welcome to openstudy....your gonna like it here :)